Molar mass of ethanol (C2​H5​OH) is: (Atomic masses: C = 12 u, H = 1 u, O = 16 u)

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
Dear student

C2H5OH :

Molecular weight of carbon is 12

Since two carbons are there 2 × 12 = 24

Molecular weight of hydrogen is 1

Since five hydrogen are there 5 × 1 = 5

Molecular weight of O and H is 16 + 1 = 17

The molar mass of ethanol  is 24 + 5 + 17 = 46